One bacterial culture is 18 days older than a second bacterial culture. In 10 days the first culture will be twice the age of the second. What is the present age of both cultures?
I know that:
The age of the bacterial culture in ten days = 2 times age of the second bacterial culture
But I don't know where to go from here.

Answer by Greenfinch(383)

younger one is x days old, older one is x + 18 days old
In ten days 2(x + 10) = x + 28
Or 2x + 20 = x + 28
x = 8
Check : now 8 days and 26 days old
in 10 days younger one is 18 days and older one is 36 days old